Sita expects her future earnings to be worth Rs. 100. If she falls ill, her expected future
earning will be Rs 25. There is a belief that she may fall ill with probability of 

while
the probability of remaining in good health is

. Let her utility function be given as U(y) =



. suppose that an insurance company offers to fully insure sita against loss of earnings
caused by illness against an actuarially fair premium.
(a) Will Sita accepts the insurance. Explain.
(b) What is the maximum amount that Sita would pay for the insurance?
1
Expert's answer
2016-12-21T09:12:11-0500
U(y) = y1/2;
(a) Sita may accept the insurance, if the probability of illness is too
high and the insurance payment is affordable. As p = 1/2, Sita will be able to
refuse the insurance.
(b) The maximum amount that Sita would pay for the insurance can be calculated according to probabilities:
i(max) = 100*1/2 + 25*1/2 = 50 + 12.5 = 62.5